The underlying asset price is $24. The risk free interest rate is 3% and the underlying asset volatility is 25%. What is the probability that the under underlying asset price will be greater than the strike price in three year's.


1
Expert's answer
2021-07-08T09:48:13-0400

Let underlying asset price be; K=24

Let force of interest be; d=0.02

i=0.03

d=In(1+i)

=In(1.03)=0.02


K=Soedt

24=Soe(0.02*3)

24=1.0618So

So=22.60

Rate of depreciation=25%=0.25

Price after 3 years=22.6*(1.25)-3

=11.5712

Pr(K>So)=1-Pr(K<=So)=1-11.5712/24

=0.5179